OpenStudy (anonymous):

Solve the equation for y. 7x – 5y = 28 A. (-6,-14) B. (-1,-7) C. (4, 10) D. (7, 9)

OpenStudy (anonymous):

http://openstudy.com/study#/updates/56317f71e4b0b6bc456cbe07 Follow those steps and replace the number for 11

OpenStudy (anonymous):

i don't get those steps...

OpenStudy (a1234):

-9y = -8x + 11 Can you think of what to do next?

OpenStudy (anonymous):

i got logged out and it changed the question on me. lol

OpenStudy (theloshua):

7x - 5y = 28 1) (-6,-14)...x = -6 and y = -14 now we sub 7(-6) - 5(-14) = 28 -42 + 70 = 28 28 = 28 (correct) --- this one is a solution 2) (-1, -7)...x = -1 and y = -7 now we sub 7(-1) - 5(-7) = 28 -7 + 35 = 28 28 = 28 (correct) -- this one works as well

OpenStudy (anonymous):

ok thanks

OpenStudy (theloshua):

yepp ;)

OpenStudy (theloshua):

medal would be appreciated